OpenStudy (anonymous):

Find the three geometric means between 1/2and 8.

Parth (parthkohli):

Geometric mean of a,b is. \(\Large \color{MidnightBlue}{\Rightarrow \sqrt{ab} }\)

Parth (parthkohli):

Find the geometric mean, then find the geometric mean of that and 1/2 and then the geometric mean of the original geometric mean and 8.

OpenStudy (anonymous):

Find the three geometric means between 1/2 and 8. Let the means be a, b, c. => b^2 = (1/2)(8) => b = 2 (It must be +ve.) => a^2 = 1 and c^2 = 16. => a = ±1, c = ±4. => means are ±1, 2, ±4.

OpenStudy (anonymous):

Thank you soo much! I was stuck on this forever.